A certain company manufactures a certain product, P , from three raw materials, combined according to a production function given by
2/5 1/5 1/5 P=50x y z .
Since the budget available for the purchase of the three raw materials of 24000 euros and knowing that they cost for unit 80, 12 and 10 euros, respectively, for x, y and z, what combination should you have regarding the purchase of products x, y and z, in order to maximize production?
1
Expert's answer
2020-05-10T12:16:08-0400

"\phi(u)=-50x^{2/5}y^{1/5}z^{1/5}\\
\min \phi(u)\\
g_0(u)=80x+12y+10z-24000\\
g_0(u)\leq 0\\
g_1(u)=-x\\
g_1(u)<0\\
g_2(u)=-y\\
g_2(u)<0\\
g_3(u)=-z\\
g_3(u)<0\\
u=(x,y,z)\in \mathbb{R^3}"

"\frac{\partial g_0^T(u)}{\partial u}=(80\ 12\ 10)\\
\frac{\partial g_1^T(u)}{\partial u}=(-1\ 0\ 0)\\
\frac{\partial g_2^T(u)}{\partial u}=(0\ -1\ 0)\\
\frac{\partial g_3^T(u)}{\partial u}=(0\ 0\ -1)\\"

Only "g_0" may be active.

"g_0\neq 0"

Hence each acceptable point is ordinary.

We will build normal (classical) Lagrange function:

"L(x,\lambda)=\phi(u)+\lambda_0 g_0(u)+\lambda_1 g_1(u)+\lambda_2 g_2(u)\\
+\lambda_3 g_3(u)=-50x^{2/5}y^{1/5}z^{1/5}+\lambda_0(80x+12y+\\
+10z-24000)-\lambda_1x-\lambda_2y-\lambda_3z."

Due to necessary condition there exist "\lambda_0,\lambda_1,\lambda_2,\lambda_3\ (\lambda_i\geq 0)" such that:

"\frac{\partial L(u,\lambda)}{\partial x}=-20x^{-3/5}y^{1/5}z^{1/5}+80\lambda_0-\lambda_1=0\\
\frac{\partial L(u,\lambda)}{\partial y}=-10x^{2/5}y^{-4/5}z^{1/5}+12\lambda_0-\lambda_2=0\\
\frac{\partial L(u,\lambda)}{\partial z}=-10x^{2/5}y^{1/5}z^{-4/5}+10\lambda_0-\lambda_3=0\\
\lambda_0(80x+12y+10z-24000)=0\\
\lambda_1=\lambda_2=\lambda_3=0."

We have the system of equations:

"-20x^{-3/5}y^{1/5}z^{1/5}+80\lambda_0=0\\
-10x^{2/5}y^{-4/5}z^{1/5}+12\lambda_0=0\\
-10x^{2/5}y^{1/5}z^{-4/5}+10\lambda_0=0\\
\lambda_0(80x+12y+10z-24000)=0"

Using CoCalc (Sage) we find the solution of this system:


"u_0=(x_0,y_0,z_0)\\
x_0=150, y_0=500, z_0=600."

Next we will check sufficient condition.

"\frac{\partial^2 L(u_0, \lambda^*)}{\partial u^2}" will be a matrix of positive elements "a_{ij}" with 3 columns and 3 rows.

"I=u^T \frac{\partial^2 L(u_0, \lambda^*)}{\partial u^2}u\\
u^T=(x\ y\ z)\\
I=a_{11}x^2+2a_{12}xy+2a_{13}xz+\\
+a_{22}y^2+2a_{23}yz+a_{33}z^2\\
\frac{\partial g_0^T(u_0)}{\partial u}u=0\\
(80\ 12\ 10)(x\ y\ z)^T=0\\
80x+12y+10z=0\text{ --- our hyperplane}.\\
I>0 \text{ on this hyperplane}."


So we have "x=150, y=500, z=600."